# Emergent active turbulence and intermittency in dense algal suspensions of Chlamydomonas reinhardtii

### v2

@@ III Fitted PDFs
We find that the PDFs [given in Figs. 3(a), 3(c), 4(a), and 4(c) in the main paper] can be fit to a form that is the sum of an exponential and a compressed exponential. Such fits are shown in Fig. 2. For example, for the WT cells, we obtain $\mathcal{P}(\bm{u}_{x})=(1.01)e^{-22|\bm{u}_{x}|^{1.73}}+(2.09)e^{-6|\bm{u}_{x}|}$ for $\bar{\rho}=0.62$.
IV Spatial density from the microscopic image
We binarize the image in Fig. 3 such that the points, where the cells are present, are shown in white (pixel value $=1$) and the background is black (pixel value $=0$). To obtain the area fraction covered by the cells in a given region, we count the number of pixels occupied by the cells there and divide it by the total area of the region. This gives us the spatial cell-density of the suspension at a given instant of time.
@@ IV Spatial density from the microscopic image
V Videos
  • V1: This experimental video shows the spatiotemporal evolution of WT cell suspensions with $\bar{\rho}=0.49$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V2: This experimental video shows the spatiotemporal evolution of WT cell suspensions with $\bar{\rho}=0.72$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V3: This experimental video shows the spatiotemporal evolution of mbo2 cell suspensions with $\bar{\rho}=0.51$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V4: This experimental video shows the spatiotemporal evolution of mbo2 cell suspensions with $\bar{\rho}=0.71$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V5: This video shows the spatiotemporal evolution of the pseudocolor plots of the density field (left) and the vorticity $\omega$ [colour-bar unit $sec^{-1}$] field (right) for the WT cells (blue panel, $\bar{\rho}=0.49$) and the mbo2 mutant (beige panel, $\bar{\rho}=0.51$).

### v3

@@ Appendix I Fitted PDFs
We find that the PDFs [given in Figs. 4(a), 4(c), 5(a), and 5(c) in the main paper] can be fit to a form of a compressed exponential. Such fits are shown in Fig. 21 and Fig. 22. For example, for the WT cells and $\bar{\rho}=0.47$, we obtain
$\mathcal{P}(\bm{u}_{x})=(0.48)e^{-(0.81)|\bm{u}_{x}|^{1.52}}$ and $\mathcal{P}(\delta\bm{u}_{\parallel}(l))=(0.44)e^{-(0.64)|\delta\bm{u}_{\parallel}(l)|^{1.76}}$.
Figure 21: Fitted PDFs of the $x$-component of the velocity (top panel) and the longitudinal velocity increments (bottom panel) for WT cells.
Figure 22: Fitted PDFs of the $x$-component of the velocity (top panel) and the longitudinal velocity increments (bottom panel) for mbo2 cells.
@@
Appendix J Spatial density from the microscopic image
We binarize the image in Fig. 23 such that the points, where the cells are present, are shown in white (pixel value =1=1) and the background is black (pixel value =0=0). To obtain the area fraction covered by the cells in a given region, we count the number of pixels occupied by the cells there and divide it by the total area of the region. This gives us the spatial cell-density of the suspension at a given instant of time.
Appendix K Delaunay triangulation of the cell-positions
Figure 24: (a) Optical image of a WT cell suspension ($\bar{\rho}=0.50$) overlaid with Delaunay triangles. (b) Optical image of a mbo2 cell suspension ($\bar{\rho}=0.50$) overlaid with Delaunay triangles. (Scale-bar, 20 microns)
Figure. 24 shows the Delaunay triangulation of the cell-positions for WT and mbo2 suspensions ($\bar{\rho}=0.50$). We calculate the mean area of the triangles $\mu_{T}$, the median $M_{T}$ and the standard deviation $\varsigma_{T}$. For the WT cells, $\mu_{T}=51.2\mu m^{2}$, $M_{T}=47.3\mu m^{2}$ and $\varsigma_{T}=19.1\mu m^{2}$. In case of the mbo2 cells, $\mu_{T}=52.0\mu m^{2}$, $M_{T}=41.5\mu m^{2}$ and $\varsigma_{T}=31.9\mu m^{2}$. Even if the mean area for both cell types are close, there are significant differences in the median and the standard deviation. This difference in the spatial distribution of the cells can be visually seen; there appears to be much larger void spaces in the mbo2 cell suspensions compared to WT cells of similar $\bar{\rho}$. Given these void spaces in mbo2 cells, we do not interpolate the velocity vectors for $\bar{\rho}<0.6$.
Appendix L Probability distribution of the spatial density $\rho$
Figure. 25 plots the PDFs of the density $\rho$ for different grid sizes. They are shown for $\bar{\rho}\simeq 0.63$ for both WT and mbo2 cells. The mbo2 cells have a wider distribution compared to WT cells; and hence, they have more pronounced regions of high and low densities. But as the overall concentration increases, this difference dies down. Figure. 26 shows the full-width at half maxima (FWHM) of the PDFs of $\rho$ versus the grid size for two different average densities. It is seen that the FWHMs are very similar at $\bar{\rho}\simeq 0.75$.
Figure 25: PDFs of the spatial density $\rho$ computed for different grid sizes of the WT cells (top panel) and mbo2 cells (bottom panel). Here, $\bar{\rho}\simeq 0.63$.
Figure 26: FWHM of the PDFs of $\rho$ for different grid sizes (a) $\bar{\rho}\simeq 0.63$ (b) $\bar{\rho}\simeq 0.75$
Appendix M Supplementary Videos
  • V1: This experimental video shows the spatiotemporal evolution of WT cell suspensions with $\bar{\rho}=0.50$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V2: This experimental video shows the spatiotemporal evolution of WT cell suspensions with $\bar{\rho}=0.74$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V3: This experimental video shows the spatiotemporal evolution of mbo2 cell suspensions with $\bar{\rho}=0.50$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V4: This experimental video shows the spatiotemporal evolution of mbo2 cell suspensions with $\bar{\rho}=0.75$. The video is captured with a Phantom Miro C110 Camera using a 20x bright field objective in an Olympus IX83 microscope.
  • V5: This video shows the spatiotemporal evolution of the pseudocolor plots of the density field for the WT cells ($\bar{\rho}=0.50$).
  • V6: This video shows the spatiotemporal evolution of the pseudocolor plots of the the vorticity $\omega$ field for the WT cells ($\bar{\rho}=0.50$). The arrows indicate the velocity vectors; these vectors are scaled by a factor of 10 from their actual displacements for better clarity.
  • V7: This video shows the spatiotemporal evolution of the velocity vectors overlaid on the images of WT cell suspensions with $\bar{\rho}=0.50$.